9. Solve the following equation for 0 € [0, 21), show work, 2 cos? 0 = 4 sin (9)
Answer #1

Given, 2- costo = 4 sin or We know, Cosaa => Cosx 1-2 sina = 1- 2 sina na |- Cos x sinta - 2 I-Coso of sinto - 2 Then 2- Cos²Case (L): Cos 0 - 2 = 0 Cos o = 2 -1 and + 1. It is not possible, because value of cose vary between Hence solution of given